OpenStudy (anonymous):

I need help finding the vertical asymptotes of Tan(2x) from -pi to +pi

OpenStudy (anonymous):

Tan2x would be the same as sin2x/cos2x, meaning that wewould have asymptotes wherever the value of cosine is 0.

OpenStudy (anonymous):

find the vertical asymptotes of Tan (2x) where the Domain is: \[-\pi \le X \le \pi\]

OpenStudy (anonymous):

oh true never thought of it that way

OpenStudy (anonymous):

ty man

OpenStudy (anonymous):

Yep yep :3